OpenStudy (anonymous):

list y intercepts and x intercepts of the following equation. 2x+7y=6

OpenStudy (anonymous):

set x=0, y=6/7 set y=0, x=3

OpenStudy (anonymous):

Put x=0 and calculate the result, thats the y intercept Put y=0, and calculate the result, thats the x intercept

OpenStudy (anonymous):

how did you get x=6/7 and y=3

OpenStudy (anonymous):

so if I set x = 0, I get 2(0)+7y=6 now solve for y if I set y = 0, I get 2x + 7(0)=6 now solve for x

OpenStudy (anonymous):

for the first one 2(0)+7y i got 7 not 6 and if you take 2x+7(0) i got 9